40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3656     Entry time: Tue Oct 5 21:22:46 2010
Author: yuta 
Type: Update 
Category: VAC 
Subject: green beam reached PSL table 

YES ! We got the green light coming out from the OMC chamber to the PSL table !

 (Koji, Kiwamu, Yuta)

Background

As a result of the work in the chambers, the green beam reached the OMC chamber yesterday.
Today's mission was to let the green beam coming out from the chambers to the PSL table.

What we did:
1. Installed the last three steering mirrors.
  The mirrors were 532 HR mirror with AR and 1deg wedge (CVI Y2-LW-1-2050-UV-45-P/AR).
  Two of them are placed to the MC chamber. Another one was to the OMC chamber

During putting the mirrors to the MC chamber, we found that a cable tower was sitting on a position exactly where we wanted to put one of the mirrors.

So we moved the tower to the very south west corner. 

 

2. Installed a periscope to the MC chamber

The function of this periscope is to lower the beam height of the green light which is risen up by another periscope in the BS chamber.

We aligned it to the green beam, so that the beam hits the center of the mirrors on the periscope.  

 

3. Aligned the optics

We aligned the green mirrors so that we can let the green beam go out from the chamber.

Actually the inside of the OMC chamber didn't look like the same as that of our optical layout.

For example there is unknown base plate, which apparently disturbs the location of our last steering mirror.

Therefore we had to change the designed position of the steering mirror.

Now the mirror is sitting near the designed position (~ 1/2 inch off), but it's fine because it doesn't clip any 1064 beam.

 
Result:

The green beam is now hitting the north wall of the PSL table.

 

Notes:

The green beam looks having some fringes, it may be caused by a multiple reflection from a TT when the green beam goes through it. We are going to check it. 

 

Next work:

- Damping of the suspended optics

- Resurrect MC and its stable lock

- Remove MCT pickoff path

- Align optics in the main path

- Recycled Michelson lock

 IMG_3627.jpg

 

ELOG V3.1.3-